OpenStudy (adelyn_helps):

What is the mean number of calls received per day? Day 1 2 3 4 5 6 7 8 Calls Received 5 3 8 7 9 9 10 9 A. 7.5 B. 9 C. 8.5 D. 7

OpenStudy (adelyn_helps):

@Rebecca

Mehek (mehek14):

\(\dfrac{5+3+8+7+9+9+10+9}{8}\)

OpenStudy (rebecca):

Add up all the "calls received" numbers

OpenStudy (rebecca):

Then divide by the amount of calls. Which is 8.

OpenStudy (adelyn_helps):

which is 1?

Mehek (mehek14):

No

OpenStudy (adelyn_helps):

8 divided by 8 is 1

Mehek (mehek14):

add all the numbers

Mehek (mehek14):

then divide by 8

OpenStudy (adelyn_helps):

68 divided by 8?

OpenStudy (adelyn_helps):

so its 8.5?

OpenStudy (adelyn_helps):

thanks guys! :)

Mehek (mehek14):

Where'd you get 68 from

Mehek (mehek14):

8.5 is wrong

OpenStudy (adelyn_helps):

how?

OpenStudy (adelyn_helps):

I added all the numbers together got 68 and then divided by 8! So its 8.5

Mehek (mehek14):

5 + 3 + 8 + 7 + 9 + 9 + 10 + 9 = 60

Mehek (mehek14):

60/8 = 7.5

OpenStudy (rebecca):

oui. yee

Mehek (mehek14):

Might want to check your answers first before giving out wrong answers

OpenStudy (rebecca):

http://prntscr.com/b6ibh5

OpenStudy (rebecca):

ik i typed it in wrong. everyone makes mistakes ^-^

OpenStudy (adelyn_helps):

so 60 divided by 8?

Mehek (mehek14):

Yes

OpenStudy (adelyn_helps):

so its 7.5

Mehek (mehek14):

yes

OpenStudy (adelyn_helps):

thanks

Mehek (mehek14):

yw